UT 22 Aug 2020 - Night #1 of Restart

OSA: D.G. Huerta

ISp: E. Solheid

Mtn Manager: R. Hansen

ISA: B. Rothberg (Tucson - secret bunker)

Special Guest star: O. Kuhn (Tucson), R. Rothberg (Tucson -secret bunker)

Sunset: 18:66 (Local time)

Night Plan

  • DONE Test Preset (closed dome)
  • Test Preset (open)
  • DONE Restart GCS
  • AGW 5/6 Checkout
  • Observe a standard star in all modes (30-45 min for 3 images). If clear, obtain the full set of 3 images per mode; if not, make a note in the object name and obtain only 1 image per mode.
    • BD+33 2642 15h52+32d56 - at start of night (until 10pm)
    • Feige 110 23h20-05d09 - 10:30pm-morning twilight
  • Dice-5 imaging offsets
    • Test Science OB
      • Sidereal
      • Non-Sidereal

Issues:

  • Data is not being propagated to /Repository on the mountain.
  • No Repository directory created for 20200822 in Tucson ( last directory is 20200721)
  • Neon lamp appears to be back on for MODS-1 - blinking MODS1 and MODS2 exposures (shifted) shows no difference
    • Red Grating only M1R - Neon lamp on (mods1r.20200822.0119.fits)
  • MODS2-Blue sequencer Issue (IT 6386) - 1st two images of night
  • New MODS Library
    • Dual Grating M1R took 4 instead of 3 QTH1 flats for the 1.2" slitflat
    • Blue slitless flats M2B - UG5 + Vflat6 = counts 41-42k at the right end of frame (exposure time 1.5seconds) - mods2b.20200822.0118--0122 - MODS1B blue slitless <39k counts w/2sec exptime
    • MODS2B - 2.4" slit Blue-only flat - QTH1 @6 - 3 seconds - counts at extreme right edge are ~45k

Night Log

Startup for Night

UT 01:23 - logging into OBS2 to run up MODS-Bino

UT 01:33 - sievesnaps completed successfully

UT 01:41 -MODS2B sievesnap shows a seequencer issue at Y=950 and Y=2140 (mods2b.20200821.0001 fits) - taking another sievesnap

- mods2b.20200821.0002 also shows the sequencer issue

UT 01:46 - taking a 3rd sievesnap on MODS2

- M2B 0003 looks OK - updating IT 6386
Closed for Clouds - Testing Closed Dome

UT 01:53 - clouds rolling in

UT 02:28 - still closed for clouds - O.Kuhn running some LBCS tests. If we stay closed longer will run MODS cals with new library.

UT 02:33 - D.G.Huerta running test presets

UT 02:38 - sky completely clouded over

UT 03:04 - D.G. Huerta having issues with IRAF on OBS-1

UT 03:10 - clouds thinning but humiditiy starting to rise

UT 03:37 - clouds thinning, may try to open - D.G.Huerta checking outside

UT 04:08 - still closed, thermo-couples on primaries in alarm

Closed Dome MODS cals

Biases

UT 04:12 - Biases 1x2

M1B 0002-0006 - 8x, 0007-0011 -4k, 0012-0016 - 3k, ,0017-0021 - 1k

M1R 0002-0006 - 8k, 0007-0011 -4k, 0012-0016 - 3k, 0017-0021 - 1k

M2B 0004-0008 - 8k, 0009-0013 - 4k, 0014-0018 - 3k, 0019-0023 - 1k

M2R 0004-0008 - 8k, 0009-0013 - 4k, 0014-0018 - 3k, 0019-0023 - 1k

-MODS2 started first (it fnished before MODS1)

-M1R taking a long time to readout, but M1R 14 - ran red fitsflush - note to self: don't run fitsflush during a BIAS script

UT 04:32 - Biases 2x2

-MODS1 - error on script abrored with errors at line2 command NIMGS1 - race condition?

-waited a bit and race condition gone

M1B 0022-0026 8k, 0027-0031 4k, 0032-0036 3k, 0037-0041 1k

M1R 0022-0026 8k, 0027-0031 4k, 0032-0036 3k, 0037-0041 1k

M2B 0024-0028 8k, 0029-0033 4k, 0034-0038 3k, 0039-0043 1k

M2R 0024-0028 8k, 0029-0033 4k, 0034-0038 3k, 0039-0043 1k

UT 04:49 - M1R taking a while to read out again for 0025

UT 12:10 - MODS2 Bin 1x1 Biases

M2B 0162-0166 8k; 0167-0171 4k; 0172-0176 3k; 0177-0181 1k

M2R 0156-0160 8k; 0161-0165 4k; 0166-0170 3k; 0171-0175 1k

UT 12:56 - MODS1 Bin 1x1 Biases

M1B 0160-0164 8k; 0165-0169 4k; 0170-0174 3k; 0175-0179 1k

M1R 0155-0159 8k; 0160-0164 4k; 0165-0169 3k; 0170-0174 1k

Imaging Flats
Dual (Dichroic)

UT 04:51 - starting MODS2 Dual Imaging Flats

M2B 0044-0048 - u_sdss; 0049-0053 - g_sdss;

-dust doughnut upper right on u_sdss and g_sdss, smaller spots on g_sdss

M2R 0044-0048 - r_sdss; 0049-0053 i_sdss; 0054-0058 z_sdss

-dust doighnut on top right r_sdss, i_sdss,

UT 05:15 - starting MODS1 Dual Imaging Flats

M1B 0042-0046 u_sdss; 0047-0051 g_sdss

M1R 0042-0046 r_sdss; 0047-0051 i_sdss; 0052-0056 z_sdss

g_sdss - large dist doughnut upper right quadrant

r_sdss - top right qudarant - looks like the bottom left qudarant of a dust dougnut

Blue (no dichroic)

UT 05:24 - M2 Blue Imaging Flats

M2B 0054-0058 u_sdss; 0059-0063 g_sdss

-dust doughnut in upper right qudarant not visible with blue only imaging - suggests its on dichroic

UT 05:43 - M1 Blue Imaging Flats

M1B 0052-0056 u_sdss; 0057-0061 g_sdss

-same large out of focus doughnut in top right quadrant as seen in dual g_sdss flat
Red Direct

UT 05:40 - M2 Red Imagng Flats

M2R 0059-063 r_sdss; 0064-0068 i_sdss; 0069-0073 z_sdss

-dust doiughnut top right gone here too, conforming its on the dichroioc

UT 06:00 - M1 Red Imaging Flats

M1R 0057-0061 r_sdss; 0062-0066 i_sdss; 0067-0071 z_sdss

Dual Grating Calibrations

UT 05:59 - M2 Dual Slitless Flats

M2B 0064-0068 - Vflat5; 0069-0073 Vfat9 UG5

M2R 0074-0078 - Vflat5

UT 06:20 - M2 Dual Arcs

M2B 0074-0076 - NeHg, XeKr, Ar

M2R 0079-0081 - NeHg, XeKr, Ar

UT 06:30 - M2 0.6" slitflat

M2B 0077-0082

M2R 0082-0087

UT 06:50 - M2 0.8" slitflat

M2B 0083-0088

M2R 0088-0093

UT 07:10 - M2 1.0" slitflat

M2B 0089-0094

M2R 0094-0099

UT 07:29 - M2 1.2 slitflat

M2B 0095-0100

M2R 0100-0105

-needed a M2 blue fitsflush for 5 exposures

UT 07:50 - M2 2.4" slitflat

M2B 0101-0106

M2R 0106-0111

UT 08:11 - M2 5" slitflat

M2B 0107-0113

M2R 0112-0114

UT 06:26 - M1 Dual Slitless Flats

M1B 0062-0066 Vflat5 ; 0067-0071 - Vflat10 UG5

M1R 0072-0076 Vflat5

UT 06:48 - M1 Dual Arcs

M1B 0072-0074

M1R 0077-0079

UT 06:58 - M1 0.6" Slitflat

M1B 0075-0080

M1R 0080-0085

UT 07:22 - M1 0.8" slitflat

M1B 0081-0086

M1R 0086-0091

UT 07:43 - M1 1.0" slitflat

M1B 0087-0092

M1R 0092-0097

-M1B needs a fitsflush, stuck at 86 - will do when idle

-ran at UT 07:59

UT 08:04 - M1 1.2" slitflat

M1B 0093-0098

M1R 0098-0103

UT 08:25 - M1 2.4" slitflat

M1B 0099-0104

M1R 0104-0110 (took 4 instead of 3 QTH1 flats)

UT 08:48 - M1 5" slitflat

M1B - 0105-0110

M1R - 0111-0113

Red Grating Calibrations

UT 08:31 - M2 Slitless Flats

M2R 0115-0119

UT 08:43 - M2 Arcs

M2R 0120-0122

UT 08:52 - M2 0.6" slitflats

M2R 0123-0128

UT 09:06 - M2 0.8" slitflat

M2R 0129-0134

UT 09:07 - Calmode command timed out after 90seconds - aborting and re-running script

UT 09:20 - M2 1.0" slitflats

M2R 0135-0140

UT 09:33 - M2 1.2" slitflats

M2R - 0141-0146

UT 09:45 - M2 2.4" slitflats

M2R 0147-0152

UT 10:00 - M2 5" slitflats

M2R 0153-0155

UT 09:08 - M1 slitless flats

M1R 0114-0118

UT 09:19 - M1 arcs

M1R 0119-0121 - (0119 is Neon lamp - came on!)

UT 09:26 - M1 0.6" slitflat

M1R 0122--0127

- UT 09:43 GUI showed shutter open on M1R after last exposure, clicked Update, now OK

UT 09:44 - M1 0.8" slitflat

M1R 0128-0133

UT 10:00 - M1 1.0" sliftlat

M1R 0134-0139

UT 10:14 - M1 1.2" slitflat

M1R 0140-0145

UT 10:28 - M1 2.4" slitflat

M1R 0146-0151

UT 10:43 - M1 5" slitflat

M1R 0152-0154

Blue Grating Calibrations

UT 10:09 - M2 Slitless Flats

M2B 0113-0122 - near saturation in the last 5 exposures

UT 10:31 - M2 Arcs

M2B 0123-0125

UT 10:42 - M2 0.6" slitflats

M2B 0125-0131

-Calmode timeout - retrying script

UT 10:58 - M2 0.8" slitflats

M2B 0132-0137

0134-0137 needs a blue fitsflush

-ran fitsflush when exposures done

UT 11:13 - M2 1.0" slitflats

M2B 0138-0143

UT 11:27 - M2 1.2" slitflats

M2B 0144-0149

UT 11:41 - M2 2.4" slitflats

M2B 0150-0155

UT 11:55 - M2 5" slitflats

M2B 0156-0161

UT 10:52 - M1 Slitless Flats

M1B 0111-0120

UT 11:14 - M1 Arcs

M1B 0121-0123

UT 11:24 - M1 0.6" slitflats

M1B 0124-0129

UT 11:26 -Another CALMode error - aborting and resending script

UT 11:41 - M1 0.8" slitflats

M1B 0130-0135

UT 11:56 - M1 1.0" slitflats

M1B 0136-0141

UT 12:11 - M1 1.2" slitflats

M1B 0142-0147

UT 12:25 - M1 2.4" slitflats

M1B 0148-0153

UT 12:39 - M1 5" slitflats

M1B 0154-0159

UT 12:43 - need to run blue fitsflush, stuck at 0149

UT 12:35 - putting MODS2 to sleep

UT 13:30 - putting MODS1 to sleep

PROPOSED PLAN OF ACTION:

  • Telescope Wake-up ~0.5hr
    • test presets (sidereal and NS)
  • Test SX AdSec out of the beam with LBC-Blue in the beam
  • LBC Wake-up - contingent on release of LBC-Blue
    • Collimation - DoHybrid, then DOFPIA (~1hr)
    • Superfocus Sequence, PA=0, 180 (1hr)
    • Rotator Center Checks (pairs of 0,180; 90, 270) ~0.5hr
    • Test Science OB ~ 1hr
      • Sidereal
      • Non-Sidereal
    • Photometric Fields and Spectrophotometric Stars (to establish ZPs) ~ 1-1.5hrs
      • Requires multiple airmass visits (clear conditions, any seeing)
      • Cycle through all filters
      • This is now a high priority because we need to define ZPs
    • Twilight Flats
  • MODS - contingent on passing functional checkout
    • Basic Checkout
      • Verify collimation
      • Verify AGW 5/6 Hotspots are centered
      • Observe a standard star in all modes (30-45 min for 3 images). If clear, obtain the full set of 3 images per mode; if not, make a note in the object name and obtain only 1 image per mode.
      • Dice-5 imaging offsets
    • Test Science OB
      • Sidereal
      • Non-Sidereal

Closed Dome Activities

  • MODS Calibrations - using new MODS OT library
    • Imaging Flats
      • DONE Dual
      • DONE Red
      • DONE Blue
    • Spectroscopy
      • DONE Dual
      • DONE Red
      • DONE Blue
    • Bias frames
      • DONE check new 2x1 and 2x2 bias from updated MODS library
      • DONE 1x1 Binned Biases

-- BarryRothberg - 17 Aug 2020
Topic revision: r7 - 22 Aug 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback